OpenStudy (anonymous):

Write standard form of equation of a circle with radius r and center h,k) R=square root of 3; (-4,3)

jimthompson5910 (jim_thompson5910):

The center of (x-h)^2 + (y-k)^2 = r^2 is (h,k). In this case, the center is (-4,3). So h = -4 and k = 3. The number r is the radius. Since the radius is sqrt(3), we know that r = sqrt(3) (x-h)^2 + (y-k)^2 = r^2 (x-(-4))^2 + (y-3)^2 = (sqrt(3))^2 (x+4)^2 + (y-3)^2 = 3 So the equation of the circle is (x+4)^2 + (y-3)^2 = 3

OpenStudy (anonymous):

use the formula (x-h)^2 + (y-3)^2 = r^2